ADNAN MUHAMMED

Grade 12,

find the solution of the equation ydx+xdy=xy(dy-dx)

find the solution of the equation ydx+xdy=xy(dy-dx)

Grade:11

1 Answers

User
6 years ago
here divide both sides by dx then we get
y+x(dy/dx)=xy(dy/dx-1)
then divide bothsides by xy
1/x+1/y(dy/dx)=dy/dx -1
dy/dx(1/y-1)= -1-1/x
dy/dx(1- 1/y)=1+1/x
dy/dx=y(x+1)/x(y-1)